# Prompt Guiding Multi-Scale Adaptive Sparse Representation-driven Network for Low-Dose CT MAR

### v1

@@ IV-A Model Design and Optimization
where $\mathcal{P}$ denotes the forward projection (FP) operation under the LDCT condition, and n denotes the noise. The linear system in Eqn. (9) is ill-posed, and specific regularization is necessary to address solution ambiguity and suppress noise magnification. A typical optimization model can be expressed as
$\mathop{\min}\limits_{\textbf{\emph{x}}}||\mathcal{P}\textbf{\emph{x}}-\textbf{\emph{y}}||_{2}^{2}+\lambda\emph{R}(\textbf{\emph{x}})$ (10)
where $\lambda$ is a regularization parameter, and $R(\cdot)$ serves as a regularizer to provide prior information about x. To facilitate the mutual enhancement of dual domain information, we further perform joint regularization and rewrite Eqn. (10) into
$\mathop{\min}\limits_{\textbf{\emph{s}},\textbf{\emph{x}}}||\mathcal{P}\textbf{\emph{x}}-\textbf{\emph{s}}||_{2}^{2}+\alpha||\textbf{\emph{s}}-\textbf{\emph{y}}||_{2}^{2}+\lambda_{1}\emph{R}_{1}(\textbf{\emph{s}})+\lambda_{2}\emph{R}_{2}(\textbf{\emph{x}})$ (11)
where $\alpha$ is a weight parameter used to balance different data fidelity terms, and s is the clean sinogram. $\lambda_{1}$ and $\lambda_{2}$ are regularization parameters. $R_{1}(\cdot)$ and $R_{2}(\cdot)$ denote the regularization terms that enforce certain desirable properties onto s and x, respectively. In fact, sinograms with more uniform profiles are easier to restore (Wang et al., 2021). Therefore, we rewrite s as
@@
where $\odot$ represents the element-wise product, and $\tilde{\textbf{\emph{y}}}$ represents the normalization coefficient, which are typically obtained through a U-Net architecture (Wang et al., 2021). By substituting Eqn. (12) into Eqn. (11), the dual domain LDMAR problem can be formulated as
$\mathop{\min}\limits_{\tilde{\textbf{\emph{s}}},\textbf{\emph{x}}}$ $||\mathcal{P}\textbf{\emph{x}}-\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}||_{2}^{2}+\alpha||\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}-\textbf{\emph{y}}||_{2}^{2}+\lambda_{1}\emph{R}_{1}(\tilde{\textbf{\emph{s}}})$
$+\lambda_{2}\emph{R}_{2}(\textbf{\emph{x}}).$ (13)
Additionally, we utilize a learnable sparsifying frame to explore the structural information of CT images (Shi and Liu, 2023; Shi et al., 2023; Shi and Lian, 2022). The optimization model can be further expressed as
$\mathop{\min}\limits_{\tilde{\textbf{\emph{s}}},\textbf{\emph{x}}}$ $||\mathcal{P}\textbf{\emph{x}}-\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}||_{2}^{2}+\alpha||\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}-\textbf{\emph{y}}||_{2}^{2}+\lambda_{1}\emph{R}_{1}(\tilde{\textbf{\emph{s}}})$
$+\lambda_{2}||{\textbf{\emph{W}}}\textbf{\emph{x}}||_{1}$ (14)
where $||\cdot||_{1}$ represents the $\ell_{1}$ norm. To address the optimization problem defined in Eqn. (14), we employ the proximal gradient technique to alternately solve $\tilde{\textbf{\emph{s}}}$ and x. At the ($t+1$)-th iteration, $\tilde{\textbf{\emph{s}}}$ and x are solved alternately as follows.
Updating $\tilde{\textbf{\emph{s}}}$: The optimization subproblem for solving $\tilde{\textbf{\emph{s}}}$ can be derived as
$\tilde{\textbf{\emph{s}}}^{(t+1)}=$ $\mathop{\arg\min}\limits_{\tilde{\textbf{\emph{s}}}}\Big{\{}||\mathcal{P}\textbf{\emph{x}}^{(t)}-\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}||_{2}^{2}+\alpha||\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}-\textbf{\emph{y}}||_{2}^{2}$
$+\lambda_{1}\emph{R}_{1}(\tilde{\textbf{\emph{s}}})\Big{\}}.$ (15)
@@
$\tilde{\textbf{\emph{s}}}^{(t+1)}=$ $\mathop{\arg\min}\limits_{\tilde{\textbf{\emph{s}}}}\Big{\{}\frac{1}{2}||\tilde{\textbf{\emph{s}}}-[\tilde{\textbf{\emph{s}}}^{(t)}-\eta_{1}\bigtriangledown\emph{f}_{1}(\tilde{\textbf{\emph{s}}}^{(t)})]||_{2}^{2}$
$+\lambda_{1}\eta_{1}\emph{R}_{1}(\tilde{\textbf{\emph{s}}})\Big{\}}$ (16)
where $\emph{f}_{1}(\tilde{\textbf{\emph{s}}}^{(t)})=||\mathcal{P}\textbf{\emph{x}}^{(t)}-\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}^{(t)}||_{2}^{2}+\alpha||\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}^{(t)}-\textbf{\emph{y}}||_{2}^{2}$, and $\eta_{1}$ is the stepsize. Then, the updating rule of $\tilde{\textbf{\emph{s}}}$ can be deduced as
@@
Updating x: Similarly, the optimization subproblem for solving x can be derived as
$\textbf{\emph{x}}^{(t+1)}=\mathop{\arg\min}\limits_{\textbf{\emph{x}}}\Big{\{}||\mathcal{P}\textbf{\emph{x}}-\tilde{\textbf{\emph{y}}}\odot\tilde{\textbf{\emph{s}}}||_{2}^{2}+\lambda_{2}||{\textbf{\emph{W}}}\textbf{\emph{x}}||_{1}\Big{\}}.$ (19)
@@
where W represents the sparsifying frames. $S_{\varepsilon}(\cdot)$ is the soft threshold function. To enhance representation capabilities of the single-scale sparsifying frame, we replace it with the proposed PMSRNet.
[…]
Methods 1/2 Dose 1/4 Dose 1/8 Dose Average
PSNR SSIM RMSE PSNR SSIM RMSE PSNR SSIM RMSE PSNR SSIM RMSE
Single-dose models
FBP 27.81 0.4455 109.22 26.34 0.3523 127.68 23.99 0.2674 169.49 26.05 0.3551 135.46
LWFSN (Zavala-Mondragon et al., 2022) 34.01 0.7177 53.80 32.83 0.6567 60.80 31.37 0.5973 71.16 32.74 0.6572 61.92
CTformer (Wang et al., 2023a) 37.12 0.9435 39.62 36.57 0.9289 41.56 35.95 0.9241 44.15 36.55 0.9322 41.78
CNN10 (Ghani and Karl, 2018) 37.44 0.9529 38.41 36.94 0.9448 40.32 36.42 0.9387 42.23 36.93 0.9455 40.32
EDCNN (Liang et al., 2020) 37.98 0.9558 35.96 37.54 0.9511 37.46 36.90 0.9447 39.92 37.47 0.9505 37.78
OctaveNet (Won et al., 2020) 38.69 0.9635 32.85 38.12 0.9588 32.85 37.46 0.9521 37.12 38.09 0.9581 34.27
REDCNN (Chen et al., 2017) 38.76 0.9644 32.57 38.28 0.9601 34.07 37.71 0.9544 35.93 38.25 0.9596 34.19
InDuDoNet (Wang et al., 2021) 40.62 0.9651 27.19 40.03 0.9658 28.17 39.44 0.9598 29.78 40.03 0.9636 28.38
DuMSRNet 41.44 0.9750 22.91 40.45 0.9712 26.41 40.06 0.9663 26.24 40.65 0.9708 25.19
Muti-doses models
BDuMSRNet 41.90 0.9766 22.45 41.18 0.9730 24.06 40.18 0.9674 26.63 41.09 0.9723 24.38
PDuMSRNet 42.26 0.9771 20.76 41.58 0.9742 22.30 40.44 0.9667 25.28 41.43 0.9727 22.78
IV-B Dual Domain Model Interpretable Network
The existing DL-based LDMAR algorithms (Zhou et al., 2022) are data-driven, leading to a lack of model interpretability. However, model interpretability is crucial for building models based on existing knowledge and facilitating subsequent theoretical analysis. How to construct a model interpretable LDMAR network is a challenge. Inspired by the remarkable achievements of deep unfolding techniques in various computer vision tasks, we transform each iterative step in Eqn. (17) and Eqn. (22) into a corresponding network module to build dual domain LDMAR networks, where the network with the prompt guiding module is dubbed as PDuMSRNet, and the network without the prompt guiding module is dubbed as DuMSRNet. Each module of PDuMSRNet corresponds to the implementation operator in the iterative algorithm. In this manner, the proposed PDuMSRNet possesses specific physical interpretability.
@@ IV-B Dual Domain Model Interpretable Network
IV-C Loss Function
@@
$\mathcal{L}_{mse}=||{\bm{m}}\odot({\bm{x}_{\rm out}}-{\bm{x}})||_{2}^{2}+\omega_{1}||({\bm{s}_{\rm out}-\bm{s}})||_{2}^{2}$ (23)
where ${\bm{m}}$, ${\bm{x}}$ and ${\bm{s}}$ represent a binary non-metal mask, the ground truth FDCT image without metallic implants and ground truth full-dose sinogram without metallic implants, respectively. $\omega_{1}$ is a weight parameter used to balance different loss components. However, the smooth nature of MSE makes it susceptible to outliers, resulting in the creation of new artifacts. To overcome this problem, consistent with [15] and [13], we adopt the multi-scale perceptual (MSP) loss using a pre-trained ResNet-50 (He et al., 2016), which can be defined as
$\mathcal{L}_{msp}=\sum_{n=1}^{N}||\Psi_{n}({\bm{m}}\odot{\bm{x}_{\rm out}})-\Psi_{n}({\bm{m}}\odot{\bm{x}})||_{2}^{2}$ (24)
where $\Psi_{n}(\cdot)$ represents ResNet-50 for the $n$-th scale. Finally, the hybrid dual domain loss function can be formulated as
$\mathcal{L}_{\rm hybrid}=\mathcal{L}_{\rm mse}+\omega_{2}\mathcal{L}_{\rm msp}$ (25)
@@
Figure 3: Visual evaluation of different approaches on LDCT images with metallic implants under the 1/2 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks. The blue arrow indicates the reconstructed structural information. The display window is [-1000, 1000] HU.
Figure 4: Visual evaluation of different approaches on LDCT images with metallic implants under the 1/4 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks. The blue arrow indicates the reconstructed structural information. The display window is [-1000, 1000] HU.
V Experiments Setting and Results
V-A Dataset
@@
Similar to previous MAR work (Yu et al., 2021), we randomly choose 1200 FDCT images from the DeepLesion dataset (Yan et al., 2018). Additionally, the metal masks are obtained from [41], encompassing 100 metal masks of varying location, shape, and size. Subsequently, we select 1000 FDCT images, and synthesize training dataset and validation dataset with 80 and 10 metal masks, respectively. The additional set of 10 metal masks is combined with an additional set of 200 CT images, generating a test dataset of 2000 CT images. The pixel sizes of the selected 10 metal masks for testing are: [2061, 890, 881, 451, 254, 124, 118, 112, 53, 35]. We simply categorize the first four masks as large metals and the latter six as small metals for performance evaluation.
@@
For the selection of simulation parameters and scan geometry, we simulate the equiangular fan-beam projection geometry based on a 120 kVp polyenergetic X-ray source. We introduce Poisson noise to the sinogram to simulate two LDCT scenarios. Specifically, we utilize incident X-ray with $1\times 10^{5}$ photons to simulate the 1/2 dose level, $5\times 10^{4}$ photons to simulate the 1/4 dose level, and $2.5\times 10^{4}$ photons to simulate the 1/8 dose level. We adjust the size of all images to $416\times 416$ pixels and the size of the sinograms to $641\times 640$ pixels.
V-B Implementation Details and Evaluation Criteria

### v3

@@ 3.1 Model Design and Optimization
where $\mathcal{P}$ denotes the forward projection (FP) operation under the LDCT condition, and n denotes the noise. The corresponding inverse problem for this linear system in Eqn. (1) is ill-posed, and specific regularization is necessary to address solution ambiguity and suppress noise magnification. A typical optimization model can be expressed as
[…]
where $\lambda$ is a regularization parameter, and $R(\cdot)$ serves as a regularizer to provide prior information about x. To facilitate the enhancement of dual domain information and joint regularization, the optimization model can be formulated
[…]
where $\alpha$ is a weight parameter used to balance different data fidelity terms, and s is the clean sinogram. In Eqn. (3), $\lambda_{1}$ and $\lambda_{2}$ are regularization parameters. $R_{1}(\cdot)$ and $R_{2}(\cdot)$ denote the regularization terms that enforce certain desirable properties onto s and x, respectively. In fact, sinograms with more uniform profiles are easier to restore (10.1007/978-3-030-87231-1_11, 10.1007/978-3-030-87231-1_11). Therefore, we rewrite s as
@@
where $\odot$ represents the element-wise product, and $\tilde{\textbf{\emph{y}}}$ represents the normalization coefficient, which are typically obtained through a U-Net architecture (10.1007/978-3-030-87231-1_11, 10.1007/978-3-030-87231-1_11). By substituting Eqn. (4), the dual domain LDMAR problem in Eqn. (3) can be rewritten as
[…]
Additionally, we utilize a learnable sparsifying frame to explore the structural information of CT images (10198706, 10198706; 10034942, 10034942; 9763013, 9763013). The optimization model can be further expressed as
[…]
where $||\cdot||_{1}$ represents the $\ell_{1}$ norm. To address the optimization problem defined in Eqn. (6), we alternately solve $\tilde{\textbf{\emph{s}}}$ and x. At the ($t+1$)-th iteration, $\tilde{\textbf{\emph{s}}}$ and x are solved alternately as follows.
Updating $\tilde{\textbf{\emph{s}}}$: The iteration for solving $\tilde{\textbf{\emph{s}}}$ can be derived as
[…]
@@
[…]
@@
Updating x: Similarly, the iteration for solving x can be derived as
[…]
@@
where W represents the trainable sparsifying frame, and $S_{\varepsilon}(\cdot)$ is the soft threshold function defined as $soft_{\varepsilon}(\mu)=sign(\mu)max(|\mu|-\varepsilon,0)$.
3.2 Dual Domain Model Interpretable Network
While existing deep learning-based LDMAR algorithms [9825711] are data-driven, they lack model interpretability, which is crucial for building models based on existing knowledge and facilitating subsequent theoretical analysis. Constructing an interpretable LDMAR network therefore remains a challenge. To address this challenge, we leverage deep unfolding techniques, which have shown remarkable success in computer vision tasks. Each iterative step in Eqn. (9) and Eqn. (14) is transformed into a corresponding network module, thus constructing dual-domain LDMAR networks. The resulting network incorporating the prompt guiding module is referred to as PDuMSRNet, while the version without it is termed DuMSRNet. Since each module in PDuMSRNet directly corresponds to an operator in the iterative algorithm, the proposed architecture possesses inherent model interpretability.
@@ 3.2 Dual Domain Model Interpretable Network
Fig. 2: Overview of our prompt guiding multi-scale adaptive sparse representation-driven network named PMSRNet (top), our prompt guiding scale-adaptive threshold generator termed as PSATG (middle) and multi-scale coefficient fusion module called MSFuM (bottom). The PMSRNet contains three components: multi-scale sparsifying frames, PSATG and MSFuM. The PSATG is designed to generate faithful thresholds with the prior of prompt information and contains five parts: 1) The shallow feature extraction module relies on a convolution layer; 2) The deep feature extraction module models images at local, regional, and global levels by parallelizing convolutional neural networks, window attention, and anchored stripe attention; 3) The feature selection module utilizes channel attention mechanism to adaptively reweight information from different feature extraction module; 4) The threshold generating module enhances the performance of PMSRNet by generating faithful thresholds using information from the feature selection module; 5) The prompt guiding module utilizes prompt information, including LDCT images with metallic implants, metal masks, and dose maps, to guide the threshold generating process. The MSFuM which is used to fuse the multi-scale representations mainly contains a cross attention and a feed-forward network to effectively integrate features from different resolutions. Structurally, the PMSRNet adopts a four-scale architecture. At each scale, the processing flow consists of a sparse coding procedure by learnable sparsifying frame, a threshold processing procedure with the prior of prompt information, and an image generation procedure by learnable sparsifying frame, which work in sequence to ultimately produce the final reconstructed image.
4 Prompt Guiding Multi-Scale Adaptive Sparse Representation-Driven Network
4.1 Architecture Principles
Multi-scale features and cross-scale complementarity are important for solving image inverse problems. [NEURIPS2022_5b288823] proposed MSANet that effectively integrates the within-scale characteristics and cross-scale complementarity for image denoising. Inspired by this, we propose a prompt guiding multi-scale sparse representation-driven network for LDMAR. In PMSRNet, we employ a learnable sparsifying frame to filter the input data for noise removal. More precisely, this process can be expressed as
[…]
where $\tilde{\textbf{\emph{x}}}$ represents the input LDCT image with metallic implants, $\bar{\textbf{{\emph{x}}}}$ represents the filtered CT image. We extend the single-scale sparsifying frame to a multi-scale version to enhance its representational capacity. The multi-scale sparsifying frames are sequential structure where information flows between different scales while remaining unchanged within each scale. To obtain multi-scale features, we set the filter sizes in the sparsifying frame network to $\times5$, $\times7$, $\times9$, and $\times11$, with corresponding output channel numbers of 25, 49, 81, and 121, respectively. Given the critical role of the threshold in the soft threshold function, we develop a prompt guiding scale-adaptive threshold generator to generate faithful thresholds.
4.2 Prompt Guiding Scale-Adaptive Threshold Generator
According to optimization theory, spatially variant thresholds are essential for effective image denoising in the deep sparsifying transform model, as noise levels vary with radiation dose level and artifact characteristics depend on metal implants. Thus, we employ a prompt learning strategy. Recently popularized for all-in-one IR tasks (promptir, promptir), prompt learning generalizes across degradation types and levels and achieves SOTA results in multi-task IR frameworks. In LDMAR, prompts derived from dose levels and metal implants generate adaptive thresholds, addressing both low-dose CT tasks under varying dose levels and MAR task for diverse metal implants.
Contextual information is crucial for both high- and low-resolution layers. High-resolution features lose details when downsampled, while low-resolution features compromise structural consistency. To overcome these issues and leverage cross-scale complementarity, we propose a prompt guiding scale-adaptive threshold generator (PSATG) that fuses prompt features with multi-scale features from local, regional, and global levels. PSATG comprises shallow and deep feature extraction modules, a feature selection module, a threshold generation module, and a prompt guiding module. The shallow feature extraction module uses a convolutional layer. The deep feature extraction module includes branches for local (LIEB), regional (RIEB), and global (GIEB) information extraction. LIEB employs a convolutional residual structure for local information, while RIEB and GIEB utilize window attention (9879380, 9879380) and anchored stripe attention (8953616, 8953616) for regional and global information, respectively. Feature maps from the three branches are processed independently and concatenated along the channel dimension. The prompt guiding module directs threshold generation using prompt information. In processing flow of PSATG, the shallow feature extraction module first processes the input to produce $^{\hat{\bm{x}}}$. The three branches then extract local, regional, and global features from $^{\hat{\bm{x}}}$. Finally, the feature selection module reweights channels from these branches to select information more conducive to LDCT image reconstruction.
The prompt guiding module employs a prompt suite comprising LDMAR images, metal masks, and dose maps to guide the feature extraction process. Specifically, the LDCT image with metallic implants is the input of DuMSRNet. The dose map is defined as a matrix in which each element is the reciprocal of the corresponding dose level value. The metal mask is obtained through a simple thresholding segmentation of LDCT image with metallic implants. These three prompts are integrated via channel concatenation. The concatenated features are then passed to two separate fully connected layers to generate a weight vector and a bias term. An element-wise multiplication of the weights with the features is performed, followed by the addition of the bias. The resulting reweighted features are subsequently fed into the threshold generation module. Notably, the PSATG employs shared parameters across all unfolding stages during training, which enhances both model efficiency and cross-stage consistency.
As illustrated in Fig. 2, we take the LDCT images with metallic implants, the metal masks and the dose maps as prompts, and perform a channel-wise concatenation operation on them to obtain the three-channel image P, i.e.,
$\textbf{P}={\rm Concat}(\tilde{\textbf{\emph{x}}},\bm{m},\bm{d}),$ (16)
where $\bm{m}$ and $\bm{d}$ denote the metal mask and dose map, respectively.
[…]
$\textbf{\emph{k}}_{1}={\rm FC}_{1}(F_{ext}(\textbf{P})),\textbf{\emph{k}}_{2}={\rm FC}_{2}(F_{ext}(\textbf{P})),$ (17)
$\textbf{\emph{b}}_{1}={\rm FC}_{3}(F_{ext}(\textbf{P})),\textbf{\emph{b}}_{2}={\rm FC}_{4}(F_{ext}(\textbf{P})),$ (18)
$\textbf{\emph{x}}_{in}={\rm FSM}(\textbf{\emph{k}}_{1}[{\rm LIEB}(\textbf{\emph{x}}_{in}),{\rm RIEB}(\textbf{\emph{x}}_{in}),{\rm GIEB}(\textbf{\emph{x}}_{in})]+\textbf{\emph{b}}_{1}),$ (19)
and
$\textbf{\emph{x}}_{out}={\rm Conv}({\rm ReLU}({\rm Conv}({\textbf{\emph{k}}_{2}}\textbf{\emph{x}}_{in}+{\textbf{\emph{b}}_{2}})))$ (20)
where FSM($\cdot$) denotes the feature selection module, and $\textbf{\emph{x}}_{1}$, $\textbf{\emph{x}}_{2}$, $\textbf{\emph{x}}_{3}$] denotes the channel-wise concatenation of $\textbf{\emph{x}}_{1}$, $\textbf{\emph{x}}_{2}$, and $\textbf{\emph{x}}_{3}$.
4.3 Multi-Scale Coefficient Fusion Module
$\left\{\begin{aligned} \tilde{\textbf{\emph{Q}}}&=\rm R(Dconv(Conv(\textbf{\emph{Q}}))\\ \tilde{\textbf{\emph{K}}}&=\rm R(Dconv(Conv(\textbf{\emph{K}}))\\ \tilde{\textbf{\emph{V}}}&=\rm R(Dconv(Conv(\textbf{\emph{V}}))\end{aligned}\right$. (21)
[…]
Here, $\rm{SoftMax(}\bullet{)}$ is the softmax function applied to re-weight the matrix multiplication. Next, we employ the reshape and convolution operations to further extract features. In summary, the cross attention of MSFuM can be expressed as
$\rm CAttention(\textbf{\emph{Q}},\textbf{\emph{K}},\textbf{\emph{V}})=\rm Conv(R(\tilde{\textbf{\emph{V}}}\textbf{\emph{C}})).$ (23)
5 Experiments Setting and Results
5.1 Loss Function
@@
[…]
where $\bm{m}$}, $\bm{x}$} and $\bm{s}$} represent a binary non-metal mask, the ground truth FDCT image without metallic implants and ground truth full-dose sinogram without metallic implants, respectively. In Eqn. (24), $\omega_{1}$ is a weight parameter used to balance different loss components. However, the smooth nature of MSE makes it susceptible to outliers, resulting in the creation of new artifacts. To overcome this problem, consistent with [9320928] and [9928347], we adopt the multi-scale perceptual (MSP) loss using a pre-trained ResNet-50 (7780459, 7780459), which can be defined as
[…]
where the procedure $\Psi_{n}(\cdot)$ represents ResNet-50 for the $n$-th scale. Finally, the hybrid dual domain loss function can be formulated as
$\mathcal{L}_{hybrid}=\mathcal{L}_{mse}+\omega_{2}\mathcal{L}_{msp}$ (26)
@@
5.2 Dataset
@@
Similar to previous MAR work (9201079, 9201079), we randomly choose 1200 FDCT images from the DeepLesion dataset (8579063, 8579063). Additionally, the metal masks are obtained from [8331163], encompassing 100 metal masks of varying location, shape, and size. Subsequently, we select 1000 FDCT images, and synthesize training dataset and validation dataset with 80 and 10 metal masks, respectively. The additional set of 10 metal masks is combined with an additional set of 200 CT images, generating a test dataset of 2000 CT images. The pixel sizes of the selected 10 metal masks for testing are: [2061, 890, 881, 451, 254, 124, 118, 112, 53, 35]. We simply categorize the first four masks as large metals and the latter six as small metals for performance evaluation.
@@
For the selection of simulation parameters and scan geometry, we simulate the equiangular fan-beam projection geometry based on a 120 kVp polyenergetic X-ray source. We simulated three low-dose CT scenarios with Poisson noise in the sinogram. Specifically, we utilize incident X-ray containing $1\times 10^{5}$ photons for 1/2 dose level, $5\times 10^{4}$ photons for 1/4 dose level, and $2.5\times 10^{4}$ photons for 1/8 dose level. We adjust the size of all images to $416\times 416$ pixels and the size of the sinograms to $641\times 640$ pixels.
[…]
Algorithms 1/2 Dose 1/4 Dose 1/8 Dose Average SD
Single-dose models
FBP 27.81/0.4455/109.22 26.34/0.3523/127.68 23.99/0.2674/169.49 26.05/0.3551/135.46 1.57/0.0727/25.21
LWFSN 34.01/0.7177/53.80 32.83/0.6567/60.80 31.37/0.5973/71.16 32.74/0.6572/61.92 1.08/0.0492/7.13
REDCNN 38.76/0.9644/32.57 38.28/0.9601/34.07 37.71/0.9544/35.93 38.25/0.9596/34.19 0.43/0.0041/1.37
CNN10 37.44/0.9529/38.41 36.94/0.9448/40.32 36.42/0.9387/42.23 36.93/0.9455/40.32 0.42/0.0058/1.56
EDCNN 37.98/0.9558/35.96 37.54/0.9511/37.46 36.90/0.9447/39.92 37.47/0.9505/37.78 0.44/0.0045/1.63
OctaveNet 38.69/0.9635/32.85 38.12/0.9588/32.85 37.46/0.9521/37.12 38.09/0.9581/34.27 0.50/0.0047/2.01
CTformer 37.12/0.9435/39.62 36.57/0.9289/41.56 35.95/0.9241/44.15 36.55/0.9322/41.78 0.48/0.0083/1.86
LIT-Former 37.34/0.9501/37.84 36.71/0.9436/40.34 35.98/0.9350/43.36 36.68/0.9429/40.51 0.56/0.0062/2.26
InDuDoNet 40.62/0.9651/27.19 40.03/0.9658/28.17 39.44/0.9598/29.78 40.03/0.9636/28.38 0.48/0.0027/1.07
CLIPDenoising 38.25/0.9619/34.13 37.70/0.9568/36.05 36.31/0.9051/41.42 37.42/0.9413/37.20 0.82/0.0257/3.09
DuMSRNet 41.44/0.9750/22.91 40.45/0.9712/26.41 39.64/0.9649/28.06 40.51/0.9704/25.79 0.74/0.0042/2.15
Muti-doses models
AirNet 36.10/0.9277/43.89 36.04/0.9344/43.97 35.46/0.9199/46.30 35.87/0.9273/44.72 0.29/0.0059/1.12
PromptIR 38.29/0.9630/36.55 37.99/0.9605/37.38 37.60/0.9573/38.55 37.96/0.9603/37.49 0.28/0.0023/0.82
BDuMSRNet 41.90/0.9766/22.45 41.18/0.9730/24.06 40.04/0.9689/27.15 41.04/0.9728/24.54 0.77/0.0031/1.93
PDuMSRNet 42.26/0.9771/20.76 41.58/0.9742/22.30 40.44/0.9667/25.28 41.43/0.9727/22.78 0.75/0.0044/1.88
[…]
Algorithms Large Metal ———— Small Metal Average SD
Single-dose models
FBP 20.35/0.2433 23.31/0.2638 25.23/0.2742 25.43/0.2757 25.65/0.2802 23.99/0.2674 2.00/0.0132
LWFSN 29.61/0.5968 30.63/0.5950 31.99/0.5972 32.23/0.5984 32.39/0.5992 31.37/0.5973 1.08/0.0014
REDCNN 33.98/0.9370 35.83/0.9486 38.79/0.9601 39.72/0.9622 40.23/0.9643 37.71/0.9544 2.41/0.0103
CNN10 32.30/0.9126 34.26/0.9294 37.72/0.9473 38.60/0.9505 39.18/0.9537 36.42/0.9387 2.67/0.0155
EDCNN 32.84/0.9187 34.75/0.9358 38.14/0.9533 39.09/0.9566 39.67/0.9592 36.90/0.9447 2.65/0.0154
OctaveNet 33.61/0.9339 35.48/0.9460 38.61/0.9579 39.52/0.9602 40.08/0.9622 37.46/0.9520 2.50/0.0107
CTformer 32.18/0.9018 34.03/0.9174 37.17/0.9320 37.93/0.9335 38.42/0.9359 35.95/0.9241 2.43/0.0129
LIT-Former 32.62/0.9122 34.34/0.9272 37.05/0.9425 37.76/0.9455 38.14/0.9477 35.98/0.9350 2.14/0.0135
InDuDoNet 33.62/0.9259 39.04/0.9635 41.32/0.9695 41.61/0.9701 41.59/0.9703 39.44/0.9599 3.06/0.0172
CLIPDenoising 33.24/0.9029 34.75/0.9027 37.25/0.9059 37.94/0.9065 38.38/0.9077 36.31/0.9051 1.98/0.0020
DuMSRNet 35.89/0.9547 38.84/0.9649 40.78/0.9678 41.28/0.9685 41.40/0.9688 39.64/0.9649 2.17/0.0057
Muti-doses models
AirNet 31.88/0.9005 33.69/0.9129 36.64/0.9262 37.35/0.9288 37.76/0.9312 35.46/0.9199 2.42/0.0125
PromptIR 32.42/0.9313 34.67/0.9496 39.07/0.9656 40.44/0.9684 41.37/0.9715 37.60/0.9573 3.64/0.0154
BDuMSRNet 36.27/0.9587 39.32/0.9691 41.26/0.9717 41.62/0.9723 41.71/0.9728 40.04/0.9689 2.07/0.0053
PDuMSRNet 37.82/0.9583 39.57/0.9663 41.39/0.9692 41.66/0.9696 41.75/0.9700 40.44/0.9667 1.53/0.0044
5.3 Implementation Details and Evaluation Criteria